OpenStudy (bee_see):

Find the basis for the given subspace of R^3, and state its dimension for the plane 3x-4y+5z=0.

OpenStudy (amistre64):

what are 2 linearly independant vectors in the plane?

OpenStudy (bee_see):

I'm having trouble understanding that concept.

OpenStudy (amistre64):

the basis for a plane consists of 2 vectors by which any of the vectors in the plane can be represented as a linear combination of those 2

OpenStudy (amistre64):

for example, in the standard xy plane, 0,1 and 1,0 form the basis from which all the other vectors can be created if x = 0,1 and y = 1,0 then all vectors can be represented as some v = kx + cy for some constants c and k

OpenStudy (amistre64):

we are given a plane, so we need 2 vectors in it that are not a scalar of each other so that they give us 2 distinct directions to play with

OpenStudy (amistre64):

spose x=0, and y=0, what is z?

OpenStudy (bee_see):

0

OpenStudy (amistre64):

and we really dont want the vector 0,0,0 to be part of the basis .. lets try some other values lets x=1, y=0, what is z?

OpenStudy (bee_see):

can be any number? i will choose 2.

OpenStudy (amistre64):

well, when we define x and y, z is predetermined for us solveing for z: z = (4y-3x)/5 if x=1 and y=0, s=-3/5

OpenStudy (amistre64):

**z=-3/5

OpenStudy (amistre64):

if x=0 and y=1, z=4/5

OpenStudy (amistre64):

but these arent what we like to see vectors as, fractions are just ugly i spose so v1 = 1,0,-3/5 v2 = 0,1,4/5 or in a more pretty view v1 = 5,0,-3 v2 = 0,5,4 unless of course they want a unit vectors

OpenStudy (amistre64):

or orthogonal vectors ...

OpenStudy (amistre64):

is this the only basis vectors we can find for it? no

OpenStudy (bee_see):

wait, if you are solving for z, shouldn't that 5 be negative?

OpenStudy (amistre64):

why are you thinking that? tell me yor reason

OpenStudy (bee_see):